OpenStudy (anonymous):

Help!!! will give medal I need to ask if I got this right. Divide. (12t^6-8t^8)Divided(2t^3) I got 6t^3-4t^5 am I right ???

OpenStudy (anonymous):

@paki

OpenStudy (anonymous):

nnnoo..its wrong.Would u like the answer or do u want me to explain it for u?

OpenStudy (anonymous):

u can give answer but yes can u explain.

OpenStudy (anonymous):

the answer would be -2t3 • (2t2 - 3)

OpenStudy (anonymous):

hmm ok thank you

OpenStudy (anonymous):

12t6 - 8t8 Simplify —————————— 2t3 1.1 Pull out like factors : 12t6 - 8t8 = -4t6 • (2t2 - 3) 1.2 Factoring: 2t2 - 3 Theory : A difference of two perfect squares, A2 - B2 can be factored into (A+B) • (A-B) Proof : (A+B) • (A-B) = A2 - AB + BA - B2 = A2 - AB + AB - B2 = A2 - B2 1.3 t6 divided by t3 = t(6 - 3) = t3 Dividing exponents : 1.4 22 divided by 21 = 2(2 - 1) = 21 = 2 and the final result would be -2t3 • (2t2 - 3)

OpenStudy (anonymous):

yw ^^

OpenStudy (anonymous):

Ya, you are right.

OpenStudy (anonymous):

He had it right to begin with, I'm not sure if you needed to do all that work for him.

OpenStudy (anonymous):

When you have a polynomial divided by a monomial, you just make two fractions and then divide them out, it's called sum of quotients for rational expressions.